What is the length of a circular arc (in metres) substended by beta if r=6 metres and beta= 30 degrees

i.pie, ii.1, iii. 3 pie, iv.2 pie


1
Expert's answer
2021-08-26T18:40:54-0400

"arc=radius\times angle\\arc=6\times \frac{\pi}{6}\\arc=\pi"